OpenStudy (anonymous):

which of the following is equal to the expression below ? (-10^-10)^4 a.(-10)^40 b.(-10)^-10 c.(-10)^-40 d.(-10)^-6??help

OpenStudy (experimentx):

c

OpenStudy (anonymous):

i was thinking c ..

OpenStudy (anonymous):

(a^b)^c = a^(bc) so it's (-10^-10)^4 = -10^-40